OpenStudy (madeline4giles852):

The Breaker's Manufacturing Company makes 80% of a particular sensor, the Cartin Company makes 15% of them, and the Flutes Company makes the other 5%. The sensors made by Breaker's have a 4% defect rate, the Cartin sensors have a 6% defect rate, and the Flutes sensors have a 9% defect rate. If a sensor is randomly selected from the general population of all sensors, what is the probability that it is defective, given that it was made by Breaker’s? 0.04 0.05 0.15 0.80 please help will medal :)

OpenStudy (perfecta):

Let's say you had 100 sensors 80% are made by Breakers, so 0.80*100 = 80 4% of the 80 are defective, 0.04*80 = 3.2 on average 3.2/80 = 0.04

OpenStudy (madeline4giles852):

thanks so much :)!

OpenStudy (perfecta):

Your welcome if you need anymore assistance on questions you might need answering I'm here too help @madeline4giles852

OpenStudy (madeline4giles852):

awesome thank you @PerfectA